OpenStudy (anonymous):

Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

I'd use elimination because all I would need to do in order to isolate a variable is add the two equations.

OpenStudy (anonymous):

2x+5y=-3 + -15x-5y=-10 _____________________ -13x=-13 x=1 2(1)+5y=-3 2+5y=-3 5y=-5 y=-1 So x equals 1 and y equals negative one which can be written as: (1,-1)